Experiment indicate that 13.6 ev energy is required to separate a hydrogen atom into a proton and electron, that is E= -13.6 ev; Find the orbital radius of the electron in hydrogen atom.
1
Expert's answer
2020-10-12T07:49:20-0400
"E=0.5U=-\frac{ke^2}{2r}\\-13.6(1.6\cdot 10^{-19})=-\frac{(9\cdot 10^9)(1.6\cdot 10^{-19})^2}{2r}\\r=5.29\cdot 10^{-11}\ m"